A girl went to school & the teacher asked how old are you? She replied I don't know, bt I know that I'm half my mother's age. The teacher asked ok how old is your mother? She said I don't know but I know that my Dad is 5 yrs older than my mother. Teacher, how old is yor dad? She said I don't know but I know that if we add our ages the answer is 100 yrs... Nw how old is the girl, her mum & dad.
Answer by mananth(16949)

A girl went to school & the teacher asked how old are you? She replied I don't know, bt I know that I'm half my mother's age. The teacher asked ok how old is your mother? She said I don't know but I know that my Dad is 5 yrs older than my mother. Teacher, how old is yor dad? She said I don't know but I know that if we add our ages the answer is 100 yrs... Nw how old is the girl, her mum & dad.
Girl's age = x ( she does not know)
half mother's age so mother is twice her age = 2x
Dad is 5 years older than mom
Dad = 2x+5 years
Sum of their ages = 100
x+2x+2x+5 =100
5x+5=100
5x=100-5
5x=95
/5
x=19
So the girl is 19 years
so Mom is 2 *19 = 38 years
Dad you find out. their sum is 100